In a circular motion question, where an object is on a spinning disk (there is no tension force) I am required to find speed, I am given mass of the object, the radius of the spinning disk, and coefficient of friction between the two, what equation and solutions are needed to find the speed?


1
Expert's answer
2021-02-28T07:33:53-0500
"ma=m\frac{v^2}{r}\\F_f=\mu N=\mu mg\\m\frac{v^2}{r}=\mu mg\\v=\sqrt{\mu rg}"